OpenStudy (anonymous):

factor x^5-2x^4+8x^2-13x+6 completely over the set of complex numbers

OpenStudy (whpalmer4):

You asked that question already. I told you how to do it. Is there something that didn't make sense to you?

OpenStudy (anonymous):

i really need to see it step by step to fully understand

OpenStudy (whpalmer4):

How many of the steps that I gave you did you complete? What are the possible rational factors of that polynomial?

OpenStudy (anonymous):

plus or minus 1,2,3,6

OpenStudy (whpalmer4):

Right. So now you need to go and plug each of those values into \[f(x) = x^5-2x^4+8x^2-13x+6\]and see which ones make \(f(x) = 0\).

OpenStudy (anonymous):

1 makes 0

OpenStudy (whpalmer4):

Let's try the first candidate, \(x = 1\): \[f(1) = (1)^5-2(1)^4+8(1)^2 -13(1) + 6 = 1-2+8-13 +6= 0\] So that means \((x-1)\) is a factor of \(f(x)\)

OpenStudy (whpalmer4):

Our next step is to divide \(f(x)\) by \((x-1)\) to get a simplified polynomial which has the same remaining factors as \(f(x)\). What do you get when you do that?

OpenStudy (anonymous):

x^4-x^3-x^2+7x+1?

OpenStudy (whpalmer4):

close. last term is incorrect (+1), should be (-6)

OpenStudy (anonymous):

ok now here is where idk what to do next

OpenStudy (whpalmer4):

You know by the RRT that the only way we could have f(x) end with +6 is by having a factor or factors that multiplied to 6, and so if we divide off a factor ending in 1, our poly must still end with 6, right?

OpenStudy (whpalmer4):

It's worth figuring out where you made the division mistake, but let's go on. Our new polynomial is \[x^4-x^3-x^2+7x-6\]It has the same set of possible factors: \(\pm 1, \pm 2, \pm 3, \pm 6\) Let's try \(x=1\) and see if \((x-1)\) is a repeated factor. \[(1)4-(1)^3-(1)^2+7(1)-6 = 1-1-1+7-6 = 0 \]That means that \((x-1)\) is still a factor of our simplified polynomial, so we divide it by \((x-1\).

OpenStudy (whpalmer4):

Uh, \((x-1)\) What does that give us for our even simpler polynomial?

OpenStudy (anonymous):

it gives us 0 and x^3-x+6

OpenStudy (whpalmer4):

Correct! Now, once again the potential factors are \(\pm 1, \pm 2, \pm 3, \pm 6\). This time, \(x=1\) is not a zero: \[(1)^3-1+6=6\]so we have to try the others.

OpenStudy (anonymous):

ok 2?

OpenStudy (whpalmer4):

yes, \(x = 2\) is a zero of the new, slimmer, more glamorous polynomial :-) So we divide it by \((x-2)\) and what do we get?

OpenStudy (anonymous):

x^2+2x-3

OpenStudy (whpalmer4):

No, that's incorrect.

OpenStudy (anonymous):

sorry x^2+2x+3

OpenStudy (anonymous):

nevermind i dont know

OpenStudy (anonymous):

i get a remainder of 12 it should be 0 if x-2 is a factor

OpenStudy (anonymous):

x-2 is not a factor it is x+2 that is a factor

OpenStudy (whpalmer4):

Ah, let's check something: our polynomial is \[x^3-x+6\] \[(2)^3-(x)+6=8-2+6 = 12\] So 2 isn't a root after all \[(-2)^3-(-2)+6=-8+2+6 = 0\] so (x+2) is the factor, not (x-2)

OpenStudy (whpalmer4):

That's why we couldn't get the division right!

OpenStudy (anonymous):

lol so now we factor right?

OpenStudy (anonymous):

well use the quadratic formula

OpenStudy (whpalmer4):

So now you have x^2-2x+3 and you can use the quadratic or complete the square or whatever to get the two remaining complex roots

OpenStudy (anonymous):

but we will only have a total of 4 roots when there should be 5 we have one neg root one pos root and 2 com roots. we should either have 4,2,or 0 pos real roots and 2 or 0 neg real roots.

OpenStudy (whpalmer4):

No, 5 roots were found: we had 1 as a root twice, remember.

OpenStudy (whpalmer4):

You can see this if you graph the curve. At a root where the multiplicity is an even number (such as x=1 here), the curve touches the x-axis and curves away without crossing. At an odd multiplicity, the curve crosses the x-axis.